OpenStudy (anonymous):

Prove that the area bounded by a tangent at P(cp, c/p) to the rectangular hyperbola xy = c² and the asymptotes of the curve is a constant (ie, not dependent on p) Help please?

OpenStudy (anonymous):

Asymptotes are the x and y axis, and the tangent at P is x + p²y = 8p Other than that, I don't know how to do it

myininaya (myininaya):

this sounds like you may need integration? do you know integration?

OpenStudy (anonymous):

Yep. I just don't know how to apply it to this question

OpenStudy (anonymous):

as x approaches infinity, the hyperbola approaches 0, but it doesn't touch, so I'm not sure what I should integrate since I only have the lower limit

myininaya (myininaya):

well that tangent line touches the curve sometime after x=0 it touches the tangent at x=cp so assume cp>0 and use as upper limit integrate tanget line - y=c^2/x from 0 to cp and see what happens

OpenStudy (anonymous):

oh! wait hold on a sec lemme draw something

OpenStudy (anonymous):

am I trying to find the yellow bit? I was trying to work out the area of the green bit and... yeah, it didn't work

myininaya (myininaya):

oh thanks for visual i like visuals

OpenStudy (anonymous):

lol :)

myininaya (myininaya):

you need to find the x-intercept of the tangent line

OpenStudy (anonymous):

alright, I think I know how to do this now. Thanks!

myininaya (myininaya):

i will do it too and we can compare answers k :)

OpenStudy (anonymous):

okay :) i'll come back when I'm done

OpenStudy (anonymous):

oh... i'm using x + p²y = 8p as the equation of the tangent... that's not right is it... I just realised that I got that from a question where c² = 16 so do I have to work out a general one?

myininaya (myininaya):

i didnt check this but you said to find tangent at (cp,c/p) right ?

myininaya (myininaya):

ok i think this is right lol

myininaya (myininaya):

are you looking at what i did rain?

OpenStudy (anonymous):

yep, trying to figure it out

OpenStudy (anonymous):

I integrated the wrong thing. I put the equation of the hyperbola instead

myininaya (myininaya):

right according to your pic we are just wanting the area under that tangent line

myininaya (myininaya):

from x=0 to x=2cp

OpenStudy (anonymous):

y=c^2 /x =c^2 x^(-1) slope m= tangent line=-c^2 x^(-2)= -c^2 /x^2 the slope at P(cp, c/p) m=-c^2 / c^2p^2=1/p^2

OpenStudy (anonymous):

ah, okay, I got the answer :) Awesome, thanks so much!!

myininaya (myininaya):

you got 2c^2?

OpenStudy (anonymous):

I haven't finished yet, on the third last line. it's just algebra, so it should be right

OpenStudy (anonymous):

okay, got the answer :D Thank you!!!

myininaya (myininaya):

on that attachemnt i started my work in the second colum n and finished upp inthe first

myininaya (myininaya):

ok cool rain :)

OpenStudy (anonymous):

since m=slope=-c^2/x^2 you can get the EQ of the line y-y1=m(x-x1) y- c/p = (-c^2/ x^2)(x-cp) y = -c^2/x +c^3 p/x^2 +c/p you can now do the area by integration here

OpenStudy (anonymous):

good luck Rain, hope you get the correct answer